# PET-DINO: Unifying Visual Cues into Grounding DINO with Prompt-Enriched Training

### 01 Apr 2026 v1

@@ 3.3 Intra-Batch Parallel Prompting
The Intra-Batch Parallel (IBP) Prompting is shown in Figure 2. A batch of images $I_{1},I_{2},\ldots,I_{N}]$, where $N$ denotes the batch size, together with their corresponding coordinates, is fed into the visual prompt generation process. This process produces all image-level visual embeddings for each category, denoted as $V_{c}^{j}$, where $c$ represents the category id and $j$ represents the sample index in the batch. For category $c$, all visual embeddings of category $c$ within the batch are grouped together. This group is denoted as $\mathcal{V}_{c}=[\ldots,V_{c}^{j},V_{c}^{j+1},\ldots]$, which contains $n$ elements, where $n\leq N$.
@@ 4.2 Implementation Details
We develop two model variants based on Swin-T [21] and Swin-L. We use BERT-base [5] as the text encoder, and adopt the checkpoints from MM-Grounding-DINO [36]. We use AdamW [22] as the optimizer with a weight decay of 1e-4. All experiments are run on 8 GPUs with a total batch size of 32. During training, only the network modules related to visual prompt routes are updated with a learning rate of 1e-4, while the image backbone and other modules are frozen. Models are trained for 12 epochs, with the learning rate reduced by a factor of 10 at epochs 8 and 11. We use a cyclical training strategy: 8 iterations of visual prompt tuning followed by 1 iteration of text prompt tuning. For Dynamic Memory-Driven Prompting, we set $M$ to 16. The hyperparameter $d$ is set to 40 for category-rich datasets such as Object365, OpenImages, and V3Det.
@@ 4.3 Protocols and Metrics
MM-Grounding-DINO [36] Text Swin-T O365, GoldG, V3Det 50.6 40.5 21.4
@@
MM-Grounding-DINO [36] Text Swin-L O365V2, OpenImage, GoldG 53.0 36.7 23.9
@@ 4.4 Comparison with Generic Detector
Text Prompt Evaluation. In this section, we evaluate the model’s performance under text prompts. Given that PET-DINO shares network components across text and visual prompt route, we examine the impact of this joint training paradigm on text prompt performance. As shown in Table 3, PET-DINO with a Swin-T backbone shows slightly lower performance compared to its predecessor, MM-Grounding-DINO [36]. This is expected due to the limited tuning of text prompt modules intended to preserve the original model’s capabilities. Crucially, despite this constraint, PET-DINO’s text prompt performance is not only well preserved but also enhanced with a more powerful backbone. Specifically, PET-DINO with a Swin-L backbone achieves improvements of 1.0 AP and 2.6 AP on COCO and LVIS minival, respectively. These results indicate that the shared network has enough capacity to accommodate the joint training of both routes. Furthermore, we speculate that training with visual prompts may facilitate the enhancement of image features for the text prompt route.
@@ 4.6 Ablation Study
Overall Ablation Study of PET-DINO. In this section, we validate the effectiveness of our visual prompt generation scheme and the two prompt-enriched training strategies. As shown in Table 6, Row 0 indicates the base model MM-Grounding-DINO, which is a text-prompted model that does not support visual prompt detection. In Row 1, the incorporation of the Alignment-Friendly Visual Prompt Generation (AFVPG) module extends MM-Grounding-DINO to support visual prompts. Although the Visual-G metric remains low at this stage, the model achieves a notable Visual-I performance of 67.0 AP. Row 2 and Row 3 incorporate Dynamic Memory-Driven Prompting (DMD) and Intra-Batch Parallel Prompting (IBP), respectively, both of which improve Visual-G performance. In particular, Row 3 substantially increases the Visual-G metric from 12.5 AP to 37.2 AP (+24.7 AP). Although Visual-I sees a slight decrease, we argue that this reflects a beneficial shift in the model’s focus: during inference, it relies less on replicating specific instances and more on recognizing the generalizable class patterns they represent. Row 4 incorporates both AFVPG and two prompt-enriched training strategies, further boosting the Visual-G metric by +3.1 AP to 40.3 AP and the Visual-I metric to 64.0 AP. Compared with Row 1, The adoption of these two prompt-enriched training strategies enables comprehensive parallel optimization, resulting in a highly efficient and unified training framework.
@@ 4.7 Analysis of Text-Prompted Pre-training
We conducted comparative experiments between training a model from scratch and inheriting a pre-trained model. The results are presented in Table 7. Our experiment demonstrates that initializing with a well-established text-prompted model as the pre-trained model yields superior performance compared to training from scratch. Specifically, the Visual-G protocol exhibits a significant performance improvement of +7.6 AP. As illustrated in Figure 3, after 12 epochs, the performance of the model trained from scratch has already stabilized, whereas inheriting the pre-trained MM-Grounding-DINO achieves a higher upper bound. This indicates that text-prompted pre-training excels at leveraging global high-level semantic representations, which helps our model understand high-level category semantics, especially in the Global-Concept Route.

### 07 Apr 2026 v2

@@ 3.3 Intra-Batch Parallel Prompting
The Intra-Batch Parallel (IBP) Prompting is shown in Figure 2. A batch of images $I_{1},I_{2},\ldots,I_{N}]$, where NN denotes the batch size, together with their corresponding coordinates, is fed into the visual prompt generation process. This process produces all image-level visual embeddings for each category, denoted as $V_{c}^{j}$, where $c$ represents the category id and $j$ represents the sample index in the batch. For category $c$, all visual embeddings of category $c$ within the batch are grouped together. This group is denoted as $\mathcal{V}_{c}=[\ldots,V_{c}^{j},V_{c}^{j+1},\ldots]$, which contains $n$ elements, where $n\leq N$.
@@ 4.2 Implementation Details
We develop two model variants based on Swin-T [21] and Swin-L. We use BERT-base [5] as the text encoder, and adopt the checkpoints from $M$-Grounding-DINO [36]. We use AdamW [22] as the optimizer with a weight decay of 1e-4. All experiments are run on 8 GPUs with a total batch size of 32. During training, only the network modules related to visual prompt routes are updated with a learning rate of 1e-4, while the image backbone and other modules are frozen. Models are trained for 12 epochs, with the learning rate reduced by a factor of 10 at epochs 8 and 11. We use a cyclical training strategy: 8 iterations of visual prompt tuning followed by 1 iteration of text prompt tuning. For Dynamic Memory-Driven Prompting, we set $M$ to 16. The hyperparameter $d$ is set to 40 for category-rich datasets such as Object365, OpenImages, and V3Det.
@@ 4.3 Protocols and Metrics
$M$-Grounding-DINO [36] Text Swin-T O365, GoldG, V3Det 50.6 40.5 21.4
@@
$M$-Grounding-DINO [36] Text Swin-L O365V2, OpenImage, GoldG 53.0 36.7 23.9
@@ 4.4 Comparison with Generic Detector
Text Prompt Evaluation. In this section, we evaluate the model’s performance under text prompts. Given that PET-DINO shares network components across text and visual prompt route, we examine the impact of this joint training paradigm on text prompt performance. As shown in Table 3, PET-DINO with a Swin-T backbone shows slightly lower performance compared to its predecessor, $M$-Grounding-DINO [36]. This is expected due to the limited tuning of text prompt modules intended to preserve the original model’s capabilities. Crucially, despite this constraint, PET-DINO’s text prompt performance is not only well preserved but also enhanced with a more powerful backbone. Specifically, PET-DINO with a Swin-L backbone achieves improvements of 1.0 AP and 2.6 AP on COCO and LVIS minival, respectively. These results indicate that the shared network has enough capacity to accommodate the joint training of both routes. Furthermore, we speculate that training with visual prompts may facilitate the enhancement of image features for the text prompt route.
@@ 4.6 Ablation Study
Overall Ablation Study of PET-DINO. In this section, we validate the effectiveness of our visual prompt generation scheme and the two prompt-enriched training strategies. As shown in Table 6, Row 0 indicates the base model $M$-Grounding-DINO, which is a text-prompted model that does not support visual prompt detection. In Row 1, the incorporation of the Alignment-Friendly Visual Prompt Generation (AFVPG) module extends $M$-Grounding-DINO to support visual prompts. Although the Visual-G metric remains low at this stage, the model achieves a notable Visual-I performance of 67.0 AP. Row 2 and Row 3 incorporate Dynamic Memory-Driven Prompting (DMD) and Intra-Batch Parallel Prompting (IBP), respectively, both of which improve Visual-G performance. In particular, Row 3 substantially increases the Visual-G metric from 12.5 AP to 37.2 AP (+24.7 AP). Although Visual-I sees a slight decrease, we argue that this reflects a beneficial shift in the model’s focus: during inference, it relies less on replicating specific instances and more on recognizing the generalizable class patterns they represent. Row 4 incorporates both AFVPG and two prompt-enriched training strategies, further boosting the Visual-G metric by +3.1 AP to 40.3 AP and the Visual-I metric to 64.0 AP. Compared with Row 1, The adoption of these two prompt-enriched training strategies enables comprehensive parallel optimization, resulting in a highly efficient and unified training framework.
@@ 4.7 Analysis of Text-Prompted Pre-training
We conducted comparative experiments between training a model from scratch and inheriting a pre-trained model. The results are presented in Table 7. Our experiment demonstrates that initializing with a well-established text-prompted model as the pre-trained model yields superior performance compared to training from scratch. Specifically, the Visual-G protocol exhibits a significant performance improvement of +7.6 AP. As illustrated in Figure 3, after 12 epochs, the performance of the model trained from scratch has already stabilized, whereas inheriting the pre-trained $M$-Grounding-DINO achieves a higher upper bound. This indicates that text-prompted pre-training excels at leveraging global high-level semantic representations, which helps our model understand high-level category semantics, especially in the Global-Concept Route.